4 Firm Ws Firm Xs Firm Ys Firm Zs Quantity Quantity Quantity Quantity Price Supplied Supplied Supplied Supplied $0 0 0 $4
Answer #1

Option 'D' is correct.

Explaination : Total quantity supplied at price $P = quantity supplied by firm w+x+y+z

Total quantity supplied at $0 : 0+0+0+0 = 0.

total quantity supplied at $4 : 2+5+4+3 = 14

total quantity supplied at $8 : 4+10+8+6 = 28

total quantity supplied at $12 : 6+15+12+9= 42

total quantity supplied at $16 : 8+20+16+12=56

total quantity supplied at $20 : 10+25+20+15 = 70

We can clearly see from above ,for every increase in price of $4,there is the increase of 14 units in total market quantity supplied.

Therefore,if the price is reduced by $4, the market quantity supplied will decrease by 14 units. Hence, Option 'D' is correct.